Home‎ > ‎Reports‎ > ‎Crew098‎ > ‎

day05


Crew 98 Log Book for January 19, 2011

Photos of the day





Commander's Check-In Report

Virgiliu Pop Reporting

Crew Physical Status: One crew member is feeling a little sick. Could be from the EVA exhaustion or from having eaten too much peanut butter.

Time Departed/Returned from EVA: 16:00-18:20

Brief Narrative of Field Mission Results: Soil samples collected.

EVA Data/Interpretations: EVA number 2; Samples to be analised.

Engineering/Hab Maintenance: Pipe leak patched.

Report Transmission Schedule: Intended by 8 PM.

Plans for Tomorrow: Will be decided in the morning.

Inventory: Yesterday we reported the apparent lack of dehydrated, ready-made meals. A visit in the attic in order to read the water level resulted in their discovery.

Also, we are missing an ATV cover, it was taken by the wind although all of them had been thoroughly secured. Mars is a windy place!

Support Requested: Pipes need to be fixed.

Miscellaneous: Second sim day.


Commander's Report

Virgiliu Pop Reporting

We all woke up at about 9 AM (we are late-risers, yet hard-workers), then we had breakfast - tortillas with peanut butter and jam (to Iulia's delight).

Haritina's morning engineering round resulted in the discovery of a grey-water pipe leak, patched by her and Dragos. Mars is indeed hostile - the plumbing situation is in dire need of being fixed. Apart from being one crew member short, the permanent resident of the hab - (the) John is not cooperating.

Haritina then discovered a stowaway - or is it martian life? - a little mouse!

Afterward we tried to fix the rover - it is not broken, but we have a power converting issue: 110 to 220 volts.

At 3 AM, the science team started to prepare for the EVA, and took the commander along in his first EVA. Mars is beautiful, yet very tiring - one can be Captain Cook, yet one cannot be a cook and go on the EVA in the same day! This was discovered once the commander returned, very tired, and was unable to 'cook' food, having to violate mission protocol and prepare some dehydrated food. I wonder if previous crews encountered the same problem!

We did not eat as yet. We are all writing our reports. The engineers did their round and at 8 pm we'll eat. We may have an early night, apart from Haritina, whose love for the stars is far greater than her exhaustion.

Engineering Report

Haritina Mogosanu Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): used

Honey (Gas Generator) (Used? Yes/No): not used

Battery Charge Level:?
29.7 V everstart
34.7 V armor plate
40.7 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator run all day OK


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.5

Propane Tank (Full/0.75/0.5/0.25/Empty):0.39

Gasoline Tank (Full/0.75/0.5/0.25/Empty):1 containers full x 5 gallons 2 half

Notes/Comments:


Oil: not checked DG said to leave it on him

Kitty Oil Quantity (# of Quarts in storage):2.5 gal

Honey Oil Quantity (# of Quarts in storage):1 gal

ATV Oil Quantity (# of Quarts in storage):on the engineering work bench
2 x 946 ml

Car Oil Quantity (# of Quarts in storage):where is it?


Water Status:

Outside Potable Water Tank Level (inches from bottom): 10

Trailer Potable Water Tank Level (inches from bottom): full

Hab Potable Water Tank Level (inches from bottom): 7.5

Potable Water Meter Reading: 410697

Notes/Comments:


GreenHab:

Condition of plants in Tank 1: plants dead please advise

Condition of plants in Tank 2: very poor only 4 plants left

Condition of plants in Tank 3: improving

Greenhab notes/ comments:


Observatory:

Telescope Used? (Yes/no) yes

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no):Yes
Oil Checked (yes/no):No
Tire Status:not checked
Odometer at end of day:-123188
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: y/n n
Oil Checked: y/n n
Fuel Consumed: filled
Tire Status:not checked
Comments: started

ATV 2 (Spirit):n
Used: y/n n
Oil Checked: y/n n
Fuel Consumed:replaced
Tire Status:not checked
Comments: started

ATV 3 (Viking I):
Used: y/n n
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status: not checked
Comments: started

ATV 4 (Viking II):
Used: y/n n
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status:not checked
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 58
Actual temperature upstairs (Farenheit): 76

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: all OK


General Engineering

1. Notes/Comments: Had a leak at the outside gray water pipe, the water was flowing through the joint.
2. One ATV cover lost to last nights wind.

Science Report (Biology)

Beatrice Gilea Reporting

Time: 16.00-18.20

Crew Members: Virgiliu Pop, Iulia Jivanescu, Beatrice Gilea

Site Location: N 38deg 24 min 19 sec
W 110 deg 47 min 00 sec

N 38deg 24min 21 ec
W 110deg 46min 51sec

Transit Mode: Pressurized rover (SUV)

Objectives: Collect rock and soil samples and bring them to the Hab for microbiological analysis

Results: samples were not yet analyzed

Lessons Learned:
It was very difficult to see and take samples because the helmets fogged;
Also it is impossible to see while the sun is oriented towards our face.
We were tired and starving when we returned from EVA

Science Report (Geology)

Iulia Jivanescu Reporting

Time: 16.00-18.20

Crew Members: Beatrice Gilea, Virgiliu Pop, Iulia Jivanescu

Site Location: N 38deg 24 min 19 sec
W 110 deg 47 min 00 sec

N 38deg 24min 21 ec
W 110deg 46min 51sec

Transit Mode: Pressurized rover (SUV)

Objectives: take rock and soil samples and analyze them for temperature, pH and radioactivity; sample them for the Biologist.
Collected rock samples in order to assess them on the Geological scale.
Took pictures to compare the Hab environment with the Martian environment

Results: samples not yet analysed

Lessons Learned:
The helmets fogged and it was difficult to take samples.
Also it is impossible to see while the sun is oriented towards our face.
We were tired and starving when we returned from EVA.


EVA 2 Report

Iulia Jivanescu Reporting

Time: 16.00-18.20

Crew Members: Beatrice Gilea, Virgiliu Pop, Iulia Jivanescu

Site Location: N 38deg 24 min 19 sec
W 110 deg 47 min 00 sec

N 38deg 24min 21 ec
W 110deg 46min 51sec

Transit Mode: Pressurized rover (SUV)

Objectives: collect samples of soil and rock

Results: samples were not yet analysed

Lessons Learned:
The helmets fogged and it was difficult to take samples.
Also, it is impossible to see while the sun is oriented towards our face.
We were tired and starving when we returned from EVA.


Journalist Report

Dragos Bratasanu Reporting

The day begun very differently today, we encountered the first problems on Mars and the engineers immediately fixed them.

The grey water pipe near the hab had a leak and it looked like it had been dripping for days. The engineers used tape and strings to solve the problem. We discovered later that we shouldn't have fixed it without checking with Mission Support and assured them that this will not happen again.

At Mission Control we still have some problems with the rover, we are not able to charge it and are looking for a fast solution.

At the end of the day, the scientists went on the second EVA, this time with the commander as the photographer in charge.

Cook's Report

Virgiliu Pop Reporting

Spreadsheet of Cook's Report.